# Good-Enough LLM Obfuscation (GELO)

### 05 Mar 2026 v1

@@ 2.2 Hybrid and Obfuscation-based Approaches
STIP’s rigid model and its critical vulnerability to open-source models make it incompatible with common deployments. PermLLM, while secure, retains the high latency of network-bound cryptographic protocols. Our protocol, GELO, is designed to overcome these specific limitations. Like PermLLM, it is secure for open-source models. However, it achieves dramatically higher performance by rejecting high-overhead MPC entirely. Instead, GELO leverages a hybrid TEE–accelerator model. By using a dynamic, per-batch secret linear transformation (an invertible matrix $A$) generated within the TEE, GELO can safely offload the vast majority of computation (the expensive matrix multiplies) to an untrusted accelerator. This approach, rooted in the computational hardness of Blind Source Separation (BSS), thwarts the statistical attacks that defeat static permutation schemes while avoiding the network latency of MPC, as will be detailed in Section 3.
@@ 3.2.1 Mitigating Gram Matrix Information Leaks
If $A$ is orthogonal (desirable for performance, as $A^{-1}=A^{\top}$), an adversary observing $U=AH$ can compute:
@@
Mitigation 1: Use a non-orthogonal $A$.
Choose $A$ to be a general invertible matrix, so that $A^{\top}A\neq I$ and the Gram matrices are masked.
@@
    Numerical stability. Orthogonal $A$ has condition number $\kappa(A)=1$. A general $A$ can have $\kappa(A)\gg 1$, amplifying accelerator-side numerical errors (e.g., BF16/FP16) in $Q=A^{-1}Y$. The TEE should generate and verify well-conditioned $A$ (e.g., enforce $\kappa(A)<100$).
@@ 4.2 Performance and Latency Analysis
Our synthetic microbenchmark uses a same-machine logical split to model the trusted/untrusted components. We obfuscate and transmit random batches between two processes running on different GPUs, rather than running end-to-end LLM inference. This setup allows us to measure the full stack overheads relevant to GELO-style offload, including $A$-generation, mixing/unmixing, and communication (IPC) latency.
@@
    For large batches ($n>2048$), overhead rises as the $\mathcal{O}(n^{3})$ cost of generating the $n\times n$ orthogonal matrix $A$ becomes the bottleneck.
@@ 4.3.4 Results: Geometric Recovery
A high error ($\gg 1.0$) indicates failure to reconstruct the structural geometry. Note that the row-side Gram matrix $G_{\text{true}}$ is not constrained by the feature-side covariance $U^{\top}U=H^{\top}H$ leaked under orthogonal $A$ (Section 3.2.1). Moreover, the attacker’s residualization (projection or subtraction) alters $U$, breaking such identities and leaving no “free” information about $G_{\text{true}}$.
@@ 4.3.5 General Blind Source Separation (BSS) Attacks
Note: Left mixing with an orthogonal $A$ preserves the feature-side covariance $U^{\top}U=H^{\top}H$, but our Gram metric is row-side and is computed after unmixing/residualization by the attacker, so this identity does not aid geometric recovery.
@@ 5.1 Why GELO works: non-identifiability by design
For a batch of $n$ token vectors, the accelerator observes the mixed hidden states $U=AH$, where $A\in\mathbb{R}^{n\times n}$ is invertible and $H\in\mathbb{R}^{n\times d}$. The observation $U$ constrains only the product $AH$. For any invertible $R\in\mathrm{GL}(n)$, the pairs $(AR^{-1},RH)$ yield the same $U$. Hence $H$ is identifiable only up to an unknown invertible transform. Without side information that ties $R$ to the true token basis, the attacker cannot uniquely recover $H$ from $U$ alone.
@@
If $A$ is freshly and independently sampled for every batch, then statistics from different batches do not align in a common coordinate system. The attacker cannot “average out” the mixing to estimate a stable inverse, in stark contrast to static obfuscation (e.g., fixed permutations), which is vulnerable to multi-run statistical attacks.
@@
ICA exploits non-Gaussianity and independence of sources under a fixed mixing matrix. Modern hidden states are high-dimensional, correlated, and structured; moreover, we refresh $A$ per batch, invalidating the fixed-mixing assumption. As a result, off-the-shelf ICA and dictionary-learning methods lack the stationary signal they need to converge.
@@ 5.2 Information-theoretic view
Here $\Sigma_{H}$ is the covariance of the hidden-state features across tokens in a batch. Without $A$, whitening reduces the problem only to an unknown orthogonal (or more generally, invertible) ambiguity; higher-order statistics that ICA would use require a fixed mixing and favorable source assumptions, which we intentionally avoid.
@@ 5.4 Batch accumulation: upper limits and empirical evidence
Fixed $A$ (not our setting).
If $A$ were fixed and the sources satisfied ICA’s identifiability assumptions (independent, suitably non-Gaussian, at most one Gaussian), then with many samples one could estimate $A$ up to permutation and scaling. GELO’s design specifically avoids this setting by refreshing $A$ each batch.
@@ 5.6 Practical implications
    If an attacker obtains enough in-batch side information to solve for the mixing matrix $A$ (e.g., many exact anchors with correct correspondences), then the remaining tokens in that batch can be de-mixed algebraically. System design should make such leakage implausible.

### 06 Mar 2026 v2

@@ 2.2 Hybrid and Obfuscation-based Approaches
STIP’s rigid model and its critical vulnerability to open-source models make it incompatible with common deployments. PermLLM, while secure, retains the high latency of network-bound cryptographic protocols. Our protocol, GELO, is designed to overcome these specific limitations. Like PermLLM, it is secure for open-source models. However, it achieves dramatically higher performance by rejecting high-overhead MPC entirely. Instead, GELO leverages a hybrid TEE–accelerator model. By using a dynamic, per-batch secret linear transformation (an invertible matrix AA) generated within the TEE, GELO can safely offload the vast majority of computation (the expensive matrix multiplies) to an untrusted accelerator. This approach, rooted in the computational hardness of Blind Source Separation (BSS), thwarts the statistical attacks that defeat static permutation schemes while avoiding the network latency of MPC, as will be detailed in Section 3.
@@ 3.2.1 Mitigating Gram Matrix Information Leaks
If AA is orthogonal (desirable for performance, as $A^{-1}=A^{\top}$), an adversary observing $U=AH$ can compute:
@@
Mitigation 1: Use a non-orthogonal AA.
Choose AA to be a general invertible matrix, so that $A^{\top}A\neq I$ and the Gram matrices are masked.
@@
    Numerical stability. Orthogonal AA has condition number $\kappa(A)=1$. A general AA can have $\kappa(A)\gg 1$, amplifying accelerator-side numerical errors (e.g., BF16/FP16) in $Q=A^{-1}Y$. The TEE should generate and verify well-conditioned AA (e.g., enforce $\kappa(A)<100$).
@@ 4.2 Performance and Latency Analysis
Our synthetic microbenchmark uses a same-machine logical split to model the trusted/untrusted components. We obfuscate and transmit random batches between two processes running on different GPUs, rather than running end-to-end LLM inference. This setup allows us to measure the full stack overheads relevant to GELO-style offload, including AA-generation, mixing/unmixing, and communication (IPC) latency.
@@
    For large batches ($n>2048$), overhead rises as the $\mathcal{O}(n^{3})$ cost of generating the $n\times n$ orthogonal matrix AA becomes the bottleneck.
@@ 4.3.4 Results: Geometric Recovery
A high error ($\gg 1.0$) indicates failure to reconstruct the structural geometry. Note that the row-side Gram matrix $G_{\text{true}}$ is not constrained by the feature-side covariance $U^{\top}U=H^{\top}H$ leaked under orthogonal AA (Section 3.2.1). Moreover, the attacker’s residualization (projection or subtraction) alters $U$, breaking such identities and leaving no “free” information about $G_{\text{true}}$.
@@ 4.3.5 General Blind Source Separation (BSS) Attacks
Note: Left mixing with an orthogonal AA preserves the feature-side covariance $U^{\top}U=H^{\top}H$, but our Gram metric is row-side and is computed after unmixing/residualization by the attacker, so this identity does not aid geometric recovery.
@@ 5.1 Why GELO works: non-identifiability by design
For a batch of $n$ token vectors, the accelerator observes the mixed hidden states $U=AH$, where $A\in\mathbb{R}^{n\times n}$ is invertible and $H\in\mathbb{R}^{n\times d}$. The observation $U$ constrains only the product AHAH. For any invertible $R\in\mathrm{GL}(n)$, the pairs $(AR^{-1},RH)$ yield the same $U$. Hence $H$ is identifiable only up to an unknown invertible transform. Without side information that ties $R$ to the true token basis, the attacker cannot uniquely recover $H$ from $U$ alone.
@@
If AA is freshly and independently sampled for every batch, then statistics from different batches do not align in a common coordinate system. The attacker cannot “average out” the mixing to estimate a stable inverse, in stark contrast to static obfuscation (e.g., fixed permutations), which is vulnerable to multi-run statistical attacks.
@@
ICA exploits non-Gaussianity and independence of sources under a fixed mixing matrix. Modern hidden states are high-dimensional, correlated, and structured; moreover, we refresh AA per batch, invalidating the fixed-mixing assumption. As a result, off-the-shelf ICA and dictionary-learning methods lack the stationary signal they need to converge.
@@ 5.2 Information-theoretic view
Here $\Sigma_{H}$ is the covariance of the hidden-state features across tokens in a batch. Without AA, whitening reduces the problem only to an unknown orthogonal (or more generally, invertible) ambiguity; higher-order statistics that ICA would use require a fixed mixing and favorable source assumptions, which we intentionally avoid.
@@ 5.4 Batch accumulation: upper limits and empirical evidence
Fixed AA (not our setting).
If AA were fixed and the sources satisfied ICA’s identifiability assumptions (independent, suitably non-Gaussian, at most one Gaussian), then with many samples one could estimate AA up to permutation and scaling. GELO’s design specifically avoids this setting by refreshing AA each batch.
@@ 5.6 Practical implications
    If an attacker obtains enough in-batch side information to solve for the mixing matrix AA (e.g., many exact anchors with correct correspondences), then the remaining tokens in that batch can be de-mixed algebraically. System design should make such leakage implausible.